### v3

Abstract
The use of solar energy has experienced explosive growth in recent years due to the comparative advantages it offers. Here, we present an innovative tool for using the light sensor of smartphones for solar irradiance experiments in particular to measure the primary component of ground-level solar radiation: the direct normal solar irradiance. We provide comprehensive guidelines for using this sensor in a variety of situations and show how to calibrate it using easy available clear-sky satellite data. In addition, as an alternative approach, we also incorporate a comparison with professional reference measurement estimates to demonstrate the consistency of the proposed methods. This experiment can be performed with the vast majority of available smartphones, some readily available materials plus an open location in clear sky conditions for a lapse of a couple of hours. In this way, the smartphone becomes a valuable tool for measuring the solar resource and to propose several experiments focusing on the physics of solar radiation and its interaction with the Earth’s atmosphere, exploring solar variations across locations, cloud conditions, and time scales.
keywords:
[…]
1 Introduction
Photovoltaic energy has experienced spectacular growth in recent years BP (2021); REN21 (2021). This growth contrasts with the little attention given to solar resources in science and engineering curricula. In general, the study of solar radiation is predominantly approached from a theoretical standpoint, potentially due to the technical complexities and high costs associated with measurement instruments. As a result, experiments involving the measurement of solar irradiance and the determination of its uncertainty are typically reserved for advanced programs or specialized laboratories. It is also worth mentioning that this subject matter provides a platform to develop various cross-cutting competencies, including measurement techniques and environmental stewardship.
The main objective of this work is to show how to calibrate a light sensor like the ones incorporated in most smartphones to measure the solar resource. The achievement of this objective allows to dispose of an instrument to measure solar radiation for other activities and/or laboratories and also provides the opportunity to learn about the solar resource in general. In particular, we show how to measure broadband direct solar irradiance at normal incidence (DNI), the main component of solar radiation at ground level. Direct irradiance is the portion of the incident radiation that arrives directly from the solar disk without being absorbed or scattered in the Earth’s atmosphere. This component is essential to evaluate the performance of concentrating solar applications and to estimate the solar irradiance available on an inclined plane, necessary for solar photovoltaic and low-temperature solar thermal applications in which it is usual practice to tilt the solar collection surfaces.
The use of smartphones as a measurement instrument, which have also shown innovative contributions to the teaching of Physics due to the availability of various built-in sensors (for recent reviews see Ref. Monteiro and Martí (2022); Gil and Di Laccio (2017)). Although the type of sensor varies the type of sensor varies considerably depending on the specific model, in general, acceleration, magnetic field or luminosity sensors are the most common. The latter, which we will use in this work, has been used in relatively few experiments. It is worth mentioning its use for experimentation with polarized waves Monteiro et al. (2017) or in problems motivated in astronomical situations Di-Laccio et al. (2021). Among the advantages of using smartphone sensors we can mention that it is possible to measure simultaneously with several sensors mon (2019) or supplementted with video analysis mon (2021). In the framework of experiences aimed at students, the determination of uncertainty and the study of the fluctuations of these sensors cannot be left aside Monteiro et al. (2021)
To reach our goal, we describe two procedures for calibrating the sensor, one based on a professional reference measurement and the other using publicly-available satellite estimates of ground level clear-sky irradiance. We compare the calibration procedures and show that both are feasible. By bridging this gap about calibration, we enable the measurements and instructional laboratory to be done with only the smartphone and manual positioning, if required. We must emphasize that the professional equipment mentioned in the paper, including tracking, is not indispensable to carry out the proposal, and is included only to show the reader that the smartphone is a calibrable and reliable instrument.
As the topics concernig the solar resource are not usually covered in physics courses, next section introduces the various magnitudes and usual geometric calculations in the area of solar energy and lighting. Section 3 shows the experimental setup carried out in this work for the calibration of the illuminance sensor of a smartphone for the measurement of DNI. This section also introduces the solar satellite estimates, which are used here as an alternative reference data set for smartphone calibration. Section 4 describes the calibration process and its uncertainty evaluation. Finally, Section 5 summarizes the conclusions.
2 Theoretical framework: solar radiation
2.1 Basis of the interaction between solar radiation and the atmosphere
The solar irradiance, , is the incident power per unit normal surface whose beam comes from the Sun. The solar irradiance at the top of the atmosphere, , varies because of two factors: a variation of about % due to the elliptical character of the Earth’s orbit around the Sun, and also, it present small variations due to oscillations in the solar activity, typically below % Fröhlich (2006). The averaged solar irradiance at the top of the atmosphere on a surface normal to the Earth-Sun direction and when the Earth is at a distance equal to the mean Earth-Sun distance (1 Astronomical Unit or AU) is known as the solar constant Kopp and Lean (2011), Wm-2. In this way, the seasonal variation of is obtained by multiplying the solar constant by the orbital factor, , accounting for the variation of the Earth-Sun distance and it can be approximated with an uncertainty of 0.25% by
where is the ordinal day number (going from 1, Jan. 1st. to 365 Dec. 31) Iqbal (1983); Duffie and Beckman (2013). The extraterrestrial irradiance at normal incidence is obtained then as .
Once solar irradiance penetrates the Earth’s atmosphere, it interacts with various atmospheric components such as air, aerosols, water vapor, and cloudiness. This interaction leads to scattering in multiple directions, with some of the irradiance being absorbed by these components, while the remaining portion is reflected back into space. The combination of these two components on a horizontal plane is known as the global horizontal irradiance (GHI), denoted as , which represents the solar energy magnitude most commonly measured on the Earth’s surface. Several methods can be employed to measure this quantity, including the use of photodiodes, calibrated photovoltaic cells, or thermopile pyranometers. Among these options, the last offers the highest precision.
Direct normal irradiance, denoted as , is less frequently measured since its continuous measurement requires fine solar tracking mechanisms that ensure the measuring equipment to be aligned at all times pointing to the solar disk. The measuring instrument, the pyroheliometer, is equipped with a collimating tube that filters any irradiance that does not come from its normal direction, with a convention aperture of 5 sr of solid angle, corresponding to a typical solar disk. The size of the solar disk observed from Earth depends on atmospheric conditions. In the presence of high humidity, for example, the perceived solar disk enlarges due to a larger size of the circumsolar region. The solid angle of 5 sr associated with the solar disk is, in effect, a convention value. The standard that classifies solar radiation measurement instruments is the ISO 9060:2018 which establishes categories according to the quality of the equipment (offset, angular error, response time, among others) and the corresponding uncertainty. After measuring the DNI, the atmospheric transmittance can be estimated as , a quantity that enables understanding the amount of available solar radiation for capture and conversion into solar energy.
It is useful to define the main angles related to solar radiation. One of them is the solar zenith angle, denoted as , which represents the angle between the direction of the Sun and the local vertical (referred to as the local zenith) as shown in Fig. 1. The cosine of this angle appears recurrently in expressions related to solar radiation, especially for magnitudes projected onto the horizontal plane, and its calculation is carried out according to Duffie and Beckman (2013)
where is the latitude, is the solar declination angle and is the hour angle shown in Fig. 1. Latitude is the angle between the Earth’s equator (parallel 0°) and the site of interest (indicated by in the figure), along the observer’s meridian. By convention, latitudes are positive north of the equator and negative south of the equator. Solar declination is the angle formed by the Earth-Sun line with the Earth’s equatorial plane, and can be calculated in radians with good precision through the expression:
Finally, the hour angle is the angle on the equatorial plane between the observer’s meridian and the solar meridian. This angle varies with the apparent position of the Sun respect to the Earth and it is calculated from the time label associated with each measurement. Figure 1 also depicts a fourth relevant angle, , the longitude of the observer measured from the Greenwich meridian (defined as °).
The hour angle is related to the solar time of the site, , according to
Indeed, this angle vanishes at solar noon ( hours), i.e. when the solar meridian coincides with the observer’s meridian, and grows at a rate of radians per hour, the speed of rotation of the Earth. To complete the calculation, it remains to link the solar local time with the local standard time, , expressed according to a given UTC time zone associated to a central meridian, . For example, the time in UTC-3 is associated with a meridian of °. The relationship between both hours includes the so-called equation of time, , and is defined by
where is the local standard time expressed in hours and fractions, and are the signed longitude in decimal degrees (negative for West longitudes and positive for East longitudes) of the site and the reference UTC, respectively, and is expressed in minutes. can be calculated as a Fourier expansion as proposed by Spencer Spencer (1971).
2.2 Attenuation in the atmosphere
The mass of air or relative optical mass, , is a dimensionless quantity that is defined as the quotient between the amount of mass of a certain component -th of the atmosphere that a beam of radiation covers in its trajectory and the one that it would cover in a vertical path, that is, in the direction of the zenith. Under the hypothesis of a non-anisotropic flat atmosphere, we can assume Iqbal (1983); Duffie and Beckman (2013). The uncertainty associated with this expression due to neglecting the terrestrial curvature and the refraction phenomena grows as the zenith angle is larger. However, the expression presents an uncertainty of about % for ° Iqbal (1983), and it is adequate for zenith angles between 0° and 70°. There were proposed more precise expressions that can be used for large zenith angles, around 80-90°, such as that of Kasten and Young Kasten and Young (1989). Here, as we do not consider measurements very early in the morning or very late at sunset the previous expression, the previous approximation results appropriate.
The Lambert-Beer-Bourger law describes the attenuation of a direct beam of radiation when passing through a medium. Its application to the direct normal irradiance in the atmosphere results in an exponential and spectrally selective attenuation
where is the direct spectral irradiance, is the spectral extraterrestrial irradiance corrected by the orbital factor, is the optical depth of the atmosphere and is the air mass. This equation can be derived from the differential version of the Lambert-Beer-Bouger modeling the atmosphere by a set of layers , so as the transmissivity can be expressed as . The total transmissivity results from the product of the layers, and therefore includes the effect of all different components. This is a regular assumption when modeling the interaction between the Sun’s radiation and the atmosphere Gueymard (2008); Oumbe et al. (2014).
Clear-sky models adopt the Lambert-Beer-Bouger law to describe the direct normal irradiance under ideal atmospheric conditions. In these conditions, the attenuating components encompass various factors such as air molecules (O2, N2, Ar), which are responsible for Rayleigh scattering, as well as water vapor, aerosols, ozone, and other minor gases. Ozone, although crucial for life on Earth due to its role in attenuating ultraviolet radiation, has a relatively minor contribution across the entire solar spectrum. A Rayleigh atmosphere refers to a pristine and dry atmospheric state where only the attenuation mechanism of molecular scattering is at play, leading to a clean and transparent atmosphere.
Numerous clear-sky models have been developed based on these concepts Gueymard (2012). Among them, the ESRA (European Solar Radiation Atlas) model Rigollier et al. (2000) strikes a favorable balance between simplicity and precision, making it suitable for implementation within the framework of an university experimental laboratory. This model operates using the concept of global optical depth, denoted as , which encompasses the entire solar spectrum. By incorporating the global optical depth of Rayleigh extinction, denoted as , we can express , where represents Linke’s turbidity which quantifies the number of clean and dry atmospheres that would need to be stacked to achieve the level of attenuation observed in the real atmosphere. Consequently, by adjusting a single parameter, , based on ground measurements it becomes feasible to construct a simple model for estimating DNI under clear-sky conditions as . Among the several ways to approximate the Rayleigh optical depth Kasten (1984); Louche et al. (1986); Kasten and Young (1989); Grenier et al. (1994); Kasten (1996), here we use the Kasten’s formulation Kasten (1996).
2.3 Illuminance
Photometry is the area of knowledge that is responsible for measuring the light perceived by the human eye. This quantity depends on the sensitivity of the human eye to different wavelengths in the visible region of the electromagnetic spectrum. Each wavelength has its relative weight in the response of the human eye depending on the lighting conditions (good or poor) in which the observer is. In typical lighting conditions, corresponding to a real situation at Sun, it is possible to relate the illuminance, , measured in lumens per unit area, lm/m2 or lx, with the spectral irradiance according to
where is the spectral response of the human eye and is the maximum luminous efficacy obtained with monochromatic illumination at nm.
Consequently, establishing a precise relationship between illuminance, denoted as , and broadband solar irradiance is not a straightforward task, as it depends on the spectral composition of solar irradiance at ground level within a specific portion of the spectrum. This composition, in turn, is influenced by the atmospheric conditions. This scenario resembles the calibration process for photovoltaic radiometers used to measure solar irradiance. These devices possess distinct spectral responses across different regions of the solar spectrum and are calibrated by comparing them to pyranometric radiometers with a flat spectral response (broadband) encompassing the entire solar spectrum. As a first approximation, these spectral differences can be disregarded, and the customary approach involves employing a constant or global calibration curve, determined under clear-sky conditions, to account for these effects ISO 9847:1992 (1992). This calibration methodology is adopted in the present study.
3 Materials and methods
The objective of this work is to demonstrate the usefulness of smartphones as an experimental measurement tool for direct normal solar irradiance. This requires mounting a tube around the smartphone light sensor, pointing it directly at the Sun, and then calibrating its measurement. In this study, the calibration of the equipment is conducted using two distinct approaches: (i) by comparing it to high-quality pyrheliometer data obtained from professional measurements, and (ii) by comparing it to estimates from sophisticated publicly available clear-sky models. Both calibration methods require clear-sky conditions to ensure consistent measurements and to mitigate any discrepancies associated with cloud movement.
Calibration method (i) demonstrates the potential of using smartphones for direct DNI measurements, as it utilizes a reference instrument of Secondary Standard quality. This reference instrument exhibits a measurement uncertainty of less than 1%, and it is calibrated with traceability to the World Primary Standard (WSG) at the World Radiation Center (WRC) in Davos, Switzerland. This calibration approach validates the use of smartphones as measurement devices for DNI, providing a robust and reliable reference for comparison.
Calibration method (ii) offers an alternative approach for calibrating smartphones in situations where terrestrial reference measurements are unavailable. This alternative method allows for the widespread use of smartphones as measurement instruments on a large scale and at a low cost. It utilizes sophisticated clear-sky models, which serve as a general calibration reference for smartphones. This approach addresses the need for smartphone-based measurements when traditional terrestrial reference measurements are not feasible.
3.1 Experimental measurements
The measurements were made at the Solar Energy Laboratory (LES) of the University of the Republic (Udelar). The experimental site of this laboratory is located in Salto department, in northwestern Uruguay, with geographic coordinates of (latitude) and (longitude), corresponding to the UTC-3 time zone.
In this experiment, the ambient light sensor of smartphone Samsung S5 were used thanks to the freely available Phyphox app University (2020). A simple diffuser is placed above the sensor, in this case, tracing paper printed in black, which prevents saturation of the recorded signal. A cylindrical tube painted black is also placed, which acts as a collimator for a large part of the diffuse irradiance, emulating the professional pyrheliometer collimator (see Fig. 3). We emphasize that the selection of the diffuser offers no difficulties since its only function is to attenuate the radiation to avoid saturation and therefore does not affect the calibration procedure. Lighting measurements with the light sensor are recorded by the Phyphox app on a minute scale. Once the smartphone is placed in the solar tracker, the collection begins keeping the device measuring during a interval from several minutes to hours. To protect the phone screen while the measurements are recorded throughout the day, a double sheet of white paper (A4) placed in front of the smartphone screen, acting as a radiation blocker to prevent the device from overheating as shown in Fig. 3.
The selection of an appropriate diffuser is important in achieving accurate measurements with a smartphone. The smartphone’s analog-digital converter incorporates internal electronics that adjust its gain based on the illuminance detected by the sensor. Consequently, if the solar radiation measurements are low (below klx, in this case), the equipment will automatically change its scale without notifying the user. Each scale is associated with a specific sensor saturation value, and this scaling behavior can result in erroneous measurements for significant periods when the measurement is in close proximity to the saturation value. Furthermore, for the specific smartphone used in this study, the upper limit of measurable illuminance in the higher range scale is klx, which represents the saturation threshold. As a point of reference, Michael et al. Michael et al. (2020) obtained a conversion constant of lx / W m-2, indicating that measuring 1000 W/m2 would not be feasible with our smartphone (approximately klx) without the inclusion of a diffuser. Hence, it is essential to regulate the attenuation of illuminance before it reaches the sensor, for two primary reasons: (i) to ensure that values can be accurately recorded without saturating the sensor, and (ii) to maintain consistent measurements within a specific range of scales at all times. This implies that the introduced attenuation by the diffuser must strike a balance, neither being too minimal nor too excessive, but rather falling within an intermediate range.
To illustrate the impact of different diffusers, we present the results obtained from two clear-sky days using two distinct types of diffusers, as depicted in Fig. 4. The diffusers employed were: (a) ordinary white paper with a surface mass density of 120 g/m2, and (b) black printed tracing paper. The graph indicates the illuminance measurements captured by the smartphone (indicated in black) with each diffuser, alongside the reference direct irradiance measurements obtained from the pyrheliometer (displayed in blue), and the clear-sky satellite estimates (depicted in red). The behavior of the measurement obtained using diffuser (a) is illustrated in the left panel of Fig. 4, where the various scale changes occurring at low illuminance levels between and klx are clearly observed, along with the corresponding saturation points on each scale. A similar behavior at low illuminance can be observed for diffuser (b) in the right panel of Fig. 4, but only for values below klx, with notable prominence during sunset. For measurements within the range of values exceeding klx, the equipment does not undergo scale changes, resulting in continuous and seemingly anomaly-free measurements facilitated by diffuser (b). It is also evident from the graph that the ordinary white paper diffuser attenuates the signal to a greater extent compared to the tracing paper diffuser (as depicted on the right-hand side -axis of both plots). This difference can be attributed to the higher reflectivity of ordinary white paper, particularly within the visible region of the solar spectrum. Therefore, based on our findings, we recommend the use of diffuser (b) in this study. Custom selections may be done for other smartphones, however, this point requires special attention.
To validate this method, we compared it with reference equipment. In this case, the signal generated by the CHP1 pyrheliometer (in mV) is recorded by a Fisher Scientific DataTaker DT85 data logger and is converted to irradiance (in W/m2) through the equipment constant. This measurement is the reference DNI measurement of the LES lab, and it is recorded continuously on a minute scale as an average of instantaneous measurements every 15 s.
The precise measurement of DNI presents some difficulties. To carry it out, a pyrheliometer is used, an instrument that consists of an array of thermocouples (pyranometer) attached to a collimator tube, and a precision solar tracking mechanism. If the equipment is aligned with a precision of less than °, the pyrheliometer is capable of measuring the DNI with an uncertainty about 1. The measurements are carried out in broadband, that is, the irradiance corresponding to wavelengths between 200-4000 nm (which includes the entire solar spectrum) is integrated into a single value. Figure 2 shows the experimental setup of this work consisting of a Kipp & Zonen Zonen (2020) CHP1 pyrheliometer (blue oval) and a Samsung S5 smartphone (yellow oval) assembled in a Solys2 precision solar tracker. The assembly of the smartphone is shown in Fig. 3, in particular, its location perpendicular to the axis of the black bars and the assembly of the small hand-made collimator tube for the light sensor.
3.2 Calibration based on high precision clear-sky estimates
As an alternative calibration method for places where a professional DNI measurement is not available, it is possible to use as reference accurate clear-sky estimates that use information from weather satellites and physically-consistent atmospheric models. This change in the reference implies a slight increase in the uncertainty in the determination of the calibration constants, since the DNI satellite estimate presents greater uncertainty than a ground reference measurement. There are sophisticated clear-sky models that integrate estimates of different atmospheric variables, either by satellite or by atmospheric reanalysis models, which can be considered as reference Gueymard (2008); Lefèvre et al. (2013) if they have been validated by terrestrial measurements with good concordance in various parts of the world.
One interesting choice is the CAMS Programme (2020) (Copernicus Atmosphere Monitoring Service) platform which provides free clear-sky estimates using one of these reference models for the entire globe, the McClear model Lefèvre et al. (2013). This model is based on sophisticated radiative transfer calculations from the libRadTran Mayer and Kylling (2005) library and its operational version takes the form of a multiple input table based on real-time information on the state of the atmosphere. In particular, this model uses information on aerosols, precipitable water column, and ozone obtained from the CAMS reanalysis database and Earth albedo estimates obtained by the MODIS low-orbit satellite. The CAMS reanalysis in turn assimilates weather satellite’s information to provide its modeled data. Its platform enables access to 1-minute (and other time scales) solar irradiance estimates (global, direct and diffuse) from this high-precision model by simply entering the latitude and longitude of the site of interest. The file header contains information on each of the solar magnitudes provided. For example, the dimensions of radiation are Wh m-2 (irradiation, energy in the time interval per unit area), which must be converted to W m-2 (average power per unit area) by the corresponding conversion depending on the time scale.
4 Practical use of the smartphone light sensor
After selecting a suitable diffuser (for example the black-printed tracing paper used in this work) we can compare the smartphone illuminance measurements with the DNI data. Here, a calibration function
will be used, where is the illuminance measured by the smartphone expressed in klx, are the DNI data expressed in W m-2, and and are two conversion constants to adjust. The calibration is performed with the two reference DNI data sets considered, the professional measurement of the pyrheliometer and the estimates of the McClear model. Minute measurements and estimates of November 24, 2021 were used at the LES experimental site, where clear-sky conditions were maintained throughout the day. Observing the plot of panel b in Fig. 4, for the adjustment we used only the data that meet klx, so that the smartphone sensor was always on the same scale of measurement. This value was chosen conservatively, in order to ensure measurements at intermediate values on the scale. The calibration constants for both cases are presented in Table 1 and the experimental fit in Fig. 5.
Calibration | Ground | McClear |
constant | measurement | estimation |
(W m-2 / klx) | 19.29 | 18.04 |
(W m-2) | 121.1 | 193.0 |
Uncertainty in (W m-2 / klx) | 0.05 | 0.05 |
Uncertainty in (W m-2) | 2.2 | 2.5 |
Relative uncertainty ( interval) in | 0.54% | 0.65% |
Relative uncertainty ( interval) in | 3.7% | 2.6% |
Figure 1 reveals that the constants and can be determined with low statistical uncertainty for each reference data set (terrestrial measurements and McClear estimates). These uncertainties have been obtained from the linear regression assuming a Gaussian distribution of fluctuations. The table presents the statistical uncertainties in each parameter, both absolute and relative respect to its value, the latter for an interval of , which approximately represents a confidence level of 95%. For this confidence level, the uncertainty in is less than 1% and of less than 4%, for both data sets.
It is interesting to compare the calibration curves. Satellite estimates exhibit deviations from the terrestrial measurements, so the calibration curve based on these data will present more uncertainty than the one obtained by comparison with ground measurements. As observed in Fig. 4, the McClear estimates for that day overestimate the direct irradiance. This leads the calibration curve obtained with this data set results above the calibration curve obtained with terrestrial measurements as observed in Fig. 5. Comparison of the McClear calibration curve with the measured DNI data reports a mean deviation of +2.1% (overestimation) and a mean square deviation of 2.6%. This uncertainty is above the measurement uncertainty of the reference instrument (1%), so it is distinguishable, but at the same time it is below the typical uncertainty of clear-sky satellite models (3-6%) Laguarda et al. (2020). This demonstrates that it is possible to perform the calibration based on satellite data of solar irradiance with a low uncertainty, enabling its use in the absence of reference measurements. Figure 6 shows the measurements obtained with the smartphone using both calibrations. As can be seen, in both cases a very good DNI measurement is achieved and completely acceptable for an instructional laboratory for a wide range of the day. The only downside is that the smartphone does not achieve a good measurement in the first and last minutes of the day where the illuminance is very low. Outside the diurnal range, i.e. when the sensor measurement is zero, the DNI measurement is affected by the non-zero offset of Eq. 8.
These experimental results reveal that the smartphone is an outstanding tool for measuring solar irradiance through its illuminance sensor, applicable in low-cost university physics laboratories. Furthermore, the measurement capacity achieved with the equipment is really good, even in comparison to commercial sensors. Two relevant questions regarding the measurement capacity of the smartphone arises here. The first is to evaluate the typical uncertainty of a smartphone sensor used to measure DNI with both calibration methods. To answer this question would require data acquisition for several consecutive days (2-3 weeks) similar to professional calibrations following international standards ISO 9847:1992 (1992). The second is to evaluate the stability of the calibration curve over time, that is, how robust is the sensor to gradual degradation. This would require a professional calibration of the sensor every 3 months for about a year. With this set of tests it is possible to technically evaluate the capacity limits of the smartphone sensor for the solar irradiance measurement. In fact, moderate-cost professional irradiance sensors are recommended to be calibrated once a year, and a similar recommendation can arise for the smartphone sensor. We recommend here to perform the smartphone calibration each time before its utilization (i.e. the first day of measurements). Similar studies can be carried out for the measurement of global irradiance in the horizontal plane, not only direct irradiance, which surely requires evaluating the non-planar angular response of the smartphone sensor.
5 Conclusion
This study reveals the remarkable potential of smartphone light sensors as effective tools for measuring direct solar irradiance and for introducing students to the fundamental aspects of the solar resource. To validate the possibility of calibrating smartphones, this experiment was conducted using professional solar measurement and tracking equipment t Once this feasibility was demonstrated, instructors could guide students to prepare their own smartphones similarly to the experiment, using tracing paper and a simple collimator tube, to measure direct illuminance in normal incidence. These measurements, along with direct solar irradiance data measured on-site or from satellite estimates, could be used to accurately calibrate the smartphones, providing an innovative tool to advance understanding of solar resources. Remarkably, this work demonstrated that calibration against both data sets can be performed without introducing a significant increase in uncertainty, resulting in highly reliable measurement capabilities suitable for instructional laboratories. It is important to conduct calibration on a clear-sky day, following the same principles applied in professional calibrations according to current ISO standards for commercial radiometers.
Our findings pave the way for the development of various low-cost instructional laboratories, both within traditional classroom settings and in outdoor environments. In the process, students must learn to develop important scientific-mathematical skills such as the recording and processing of experimental data, assess their quality with or without data filtering, and assigning uncertainty to simplified models for direct solar irradiance estimation at ground level. By leveraging smartphone technology, students can gain practical insights into solar irradiance measurement, fostering a deeper understanding of this important aspect of renewable energy resources.
Conceptualization, J.L.D.L. and R.A.S.; methodology, J.L.D.L. and R.A.S.; software, A.M. ; validation, J.L.D.L. and R.A.S and M.M.; data curation, J.L.D.L. and M.M.; writing—original draft preparation, J.L.D.L.; R.A.S and A.C.M.; writing—review and editing, J.L.D.L.; R.A.S and A.C.M.; visualization, J.L.D.L. and M.M. All authors have read and agreed to the published version of the manuscript.
This research received no external funding.
Not applicable.
Not applicable.
The data that support the findings of this study are available on reasonable request.
Acknowledgements.
[…]
DOAJ | Directory of open access journals |
TLA | Three letter acronym |
LD | Linear dichroism |